OpenStudy (anonymous):

]Which fraction falls between three-tenths and three-fifths? five-tenths eight-tenths four-fifths two-tenths

OpenStudy (johnweldon1993):

Well we can look at all these as a decimal three tenths = 3/10 = .30 three fifths = 3/5 = .60 so (following that way of thinking about it) which of your options fall between .30 and .60?

OpenStudy (anonymous):

eight and ten

OpenStudy (campbell_st):

well use equivalent fractions for 3/5 = 6/10 so pick the fraction between 3/10 and 6/10

OpenStudy (anonymous):

type your reply

OpenStudy (johnweldon1993):

Ahh....yes @campbell_st way of looking at it is much easier!

OpenStudy (anonymous):

thats not on there

OpenStudy (johnweldon1993):

Well I can finish from my decimal point of view I suppose... so like I said we're looking for which choice is between .30 and .60 Lets look at your last answer choice two tenths = 2/10 = .20 That's not between those 2... follow that and tell me which choice falls between .30 and .60